# Compatible Relative Lefschetz Fibrations on Admissible Relative Stein Pairs

### v2

@@ 1. Introduction
A compatible Lefschetz fibration on a Weinstein domain $W$ is a smooth map $\pi:W^{2n}\rightarrow D^{2}$ which satisfies the following conditions:
@@
    The map $\pi$ has finitely many nondegenerate distinct critical values $s_{1},...,s_{k}\in D^{2}$, and there exists an unique critical point $r_{j}\in\pi^{-1}(s_{j})$ for each $j=1,...,k$,
@@
    Near each critical point $r_{j}$ and the corresponding critical value $s_{j}$, there are local complex coordinate charts matching with the orientations of $W$ and $D^{2}$ such that the map $\pi$ locally has the form $\pi(z_{1},...,z_{n})=z_{1}^{2}+...+z_{n}^{2}$,
@@
    The restriction of the map $\pi$ to $W\backslash\pi^{-1}(\{s_{1},...,s_{k}\})$ is a locally trivial fibration whose fibers are $(2n-2)$-dimensional Weinstein subdomains of $W$.
@@
A pair $(W^{6},X^{4})$ is a relative Stein pair if $W$ is a Stein domain, $X$ is a Stein subdomain of $W$, and they satisfy the following relative handle decomposition conditions:
@@
The main idea of this definition is the following: While we create the Stein domain $W$ handle by handle, the Stein domain $X$ is evolving as a subdomain of $W$ in the meantime. Moreover, the advantage of the conditions in Definition 1.1 is that “admissible Stein pairs” (see Definition 1.8), which form a subclass of relative Stein pairs, can be described by means of special diagrams which we call relative Stein diagrams. Such diagrams are, indeed, relative versions of standard Stein diagrams of Stein surfaces introduced by Gompf in [11].
@@
where $\widetilde{W}$ is a subcritical Stein $6$-domain which splits as $\widetilde{W}=\widetilde{X}\times D^{2}$, and for each $j=1,...,\ell$ the pair $(H_{j},h_{j})$ is a relative Stein $3$-handle attached along a suspendible Legendrian $2$-link.
@@
Suspendibility of $2$-links will be introduced in Definition 3.4 and Definition 3.6. For an admissible Stein pair $(W,X)$, its subcritical part $(\widetilde{W},\widetilde{X})$ satisfies the conditions $(0)$, $(1)$ and $(2)$ of Definition 1.1 (see the proof of Proposition 4.3 or Remark 3.5). Also by definition of admissibility, the condition $(3)$ of Definition 1.1 is obviously satisfied. Hence, after attaching relative Stein $3$-handles to the relative Stein pair $(\widetilde{W},\widetilde{X})$ we obtain a new relative Stein pair $(W,X)$. Therefore, admissible Stein pairs, indeed, form a special subclass of relative Stein pairs.
@@ 2. BRIESKORN VARIETIES AS SEIFERT FILLINGS
The map $\widetilde{\pi}$ is injective on the set of its $(p-1)(q-1)$ nondegenerate distinct critical points, so it has $(p-1)(q-1)$ vanishing cycles corresponding to its critical values. The regular fiber $\widetilde{F}=\widetilde{\pi}^{-1}(\epsilon)$ of $\widetilde{\pi}$ is the Brieskorn variety $V_{\epsilon}(p,q,2)$. Note that $\pi$ and $\widetilde{\pi}$ have the same morsification term and if we intersect the fiber $\widetilde{F}$ with $z_{2}=0$ plane, we obtain the fiber $F$. Also, we have $F\setminus\partial F\subset\widetilde{F}\setminus\partial\widetilde{F}$ and $\partial F=\partial\widetilde{F}\cap F$ which means that the regular fibers of $\pi$ is properly embedded in those of $\widetilde{\pi}$. Moreover, as discussed above the critical points of $\widetilde{\pi}$ are the same as those of $\pi$ (with additional $z_{2}$-coordinates). In conclusion, the vanishing cycles in the fiber $\widetilde{F}$ of $\widetilde{\pi}$ are the suspensions of those in the fiber $F$ of $\pi$. Hence, introducing a $z_{2}^{2}$ term extends (“suspends”) the Lefschetz fibration $\pi$ on $D^{4}$ to the Lefschetz fibration $\widetilde{\pi}$ on $D^{6}$, and the monodromy $\widetilde{\varphi}_{p,q}$ of $\widetilde{\pi}$ is the monodromy of the fibered Brieskorn manifold $\Sigma(p,q,2)\subset S^{5}$ which is the product of $(p-1)(q-1)$ right-handed Dehn twists along Lagrangian spheres which are the suspensions of the corresponding vanishing cycles of $\pi$. In particular, note that the binding $\Sigma(p,q)$ of the induced open book on $S^{3}$ is properly embedded in the binding $\Sigma(p,q,2)$ of the induced open book on $S^{5}$.
@@
we see that $(-1/27,0)$ and $(1/27,0)$ are the critical points of $\pi$. By substituting these into $\pi$, we get the corresponding critical values $-0.00020322105$ and $-0.00010161052$, respectively. Then one needs to calculate the determinant of the Hessian matrix at the critical points. The Hessian matrix is given by (see [7] for a complete discussion):
@@
Refer to caption Figure 4. The regular fiber and the monodromy curves of the torus knot $\Sigma(3,2)$ (trefoil).
@@ 3. Suspended Legendrian 2-links in the complement of Brieskorn manifolds
By [18], we may assume there exists a torus knot $\Sigma(p,q)$ such that the equatorial Legendrian $1$-link $K\subset(S^{3},\xi^{3}_{std})$ is embedded in the interior of a Seifert surface $F:=V_{1}(p,q)$ of $\Sigma(p,q)$. (For simplicity, we will take $\epsilon=1$.) Also, since $F$ is the page of a compatible open book (with binding $\Sigma(p,q)$) on $(S^{3},\xi^{3}_{std})$, the framing of each component of $K$ in $F$ is equal to its contact framing in $(S^{3},\xi^{3}_{std})$ (see [1]). Therefore, (ii) and (iv) follow. By assumption, for each $j=1,...,\ell$, there exists a Lagrangian disk $D_{j}$ with an embedding $\Pi_{j}:D^{2}\rightarrow D_{j}\subset D^{4}\subset\mathbb{C}^{2}$ given by
$\Pi_{j}(r,e^{i\theta})=(x^{j}_{0}(r,\theta),y^{j}_{0}(r,\theta),x^{j}_{1}(r,\theta),y^{j}_{1}(r,\theta))$
such that the image of the map $\pi_{j}:=\Pi_{j}|_{S^{1}=\partial D^{2}}:S^{1}\hookrightarrow F\subset S^{3}$ is the embedded Legendrian
$K_{j}=\{(x^{j}_{0}(\theta),y^{j}_{0}(\theta),x^{j}_{1}(\theta),y^{j}_{1}(\theta)):=(x^{j}_{0}(1,\theta),y^{j}_{0}(1,\theta),x^{j}_{1}(1,\theta),y^{j}_{1}(1,\theta))\,|\,\theta\in S^{1}\}$
in $S^{3}$. Here, $x$-coordinates are the real parts and $y$-coordinates are the imaginary parts of the complex coordinates $(z_{0},z_{1})$ on $\mathbb{C}^{2}$. In what follows, while we construct the Legendrian $2$-knot $L_{j}$ in the Brieskorn variety $V_{1}(p,q,2)$ by suspending its equator $K_{j}$, we will also make use of the fact that $V_{1}(p,q,2)$ is obtained by suspending the Seifert surface $F=V_{1}(p,q)$ of $K$ (as observed in the proof of Proposition 2.1). Consider the Seifert surface $F\subset\mathbb{C}^{3}$ as
$F=\{(z_{0},z_{1},z_{2})\;|\;z_{0}^{p}+z_{1}^{q}=1,\;z_{2}=0\}\subset\mathbb{C}^{3}$
and so we have $K_{j}=\{(x^{j}_{0}(\theta),y^{j}_{0}(\theta),x^{j}_{1}(\theta),y^{j}_{1}(\theta),0,0)\}\subset\mathbb{C}^{3}$. Note that the coordinate functions of $\pi_{j}$ together satisfy the equation $z_{0}^{p}+z_{1}^{q}=1$ defining $F$. We’ll consider the suspension of $K_{j}$ as the union of the images of the maps $\Delta_{j}:D^{2}\times[0,1]\rightarrow\mathbb{C}^{3}$, $\widetilde{\Delta}_{j}:D^{2}\times[-1,0]\rightarrow\mathbb{C}^{3}$ given by
$\Delta_{j}(r,\theta,t)=(f(t)^{1/p}\,x^{j}_{0}(r,\theta),f(t)^{1/p}\,y^{j}_{0}(r,\theta),f(t)^{1/q}\,x^{j}_{1}(r,\theta),f(t)^{1/q}\,y^{j}_{1}(r,\theta),g(t),0)$,
$\widetilde{\Delta}_{j}(r,\theta,t)=(f(t)^{1/p}\,x^{j}_{0}(r,\theta),f(t)^{1/p}\,y^{j}_{0}(r,\theta),f(t)^{1/q}\,x^{j}_{1}(r,\theta),f(t)^{1/q}\,y^{j}_{1}(r,\theta),-g(t),0)$
where the smooth maps $f(t)$ and $g(t)$ satisfy the following properties (see Figure 5):
  • (P1)
    $f(t)+g(t)^{2}=1$,
  • (P2)
    $f(0)=1,g(0)=0$ and $g(1)=1,f(1)=0$,
  • (P3)
    On $10<t<$ 1, $f^{\prime}(t)<0$ and $g^{\prime}(t)>0$.
Figure 5. Graphs of smooth functions $f$ and $g$.
If we restrict the maps $\Delta_{j}$ and $\widetilde{\Delta}_{j}$ to $\partial D^{2}$, we see that the first four coordinates are rescaled components of $K_{j}$. The coordinates $z^{j}_{0}$ and $z^{j}_{1}$ of $K_{j}$, they together satisfy $(z^{j}_{0})^{p}+(z^{j}_{1})^{q}=1$, so by the property (P1) the image of $\Delta_{j}$ (resp. $\widetilde{\Delta}_{j}$) lies in $V_{1}(p,q,2)$:
$(f(t)^{1/p}\,z^{j}_{0}(\theta))^{p}+(f(t)^{1/q}\,z^{j}_{1}(\theta))^{q}+(g(t))^{2}=f(t)((z^{j}_{0})^{p}+(z^{j}_{1})^{q})+(g(t))^{2}=1.$
Note that $\Delta_{j}(S^{1}\times\{t\})$ and $\widetilde{\Delta}_{j}(S^{1}\times\{t\})$ are circles for each $t\in(0,1)$ and $t\in(-1,0)$, respectively, and we have
$\Delta_{j}(S^{1}\times\{0\})=K_{j}$  and  $\Delta_{j}(S^{1}\times\{1\})=(0,0,0,0,1,0)$,
$\widetilde{\Delta}_{j}(S^{1}\times\{0\})=K_{j}$  and  $\widetilde{\Delta}_{j}(S^{1}\times\{-1\})=(0,0,0,0,-1,0)$.
Thus, after smoothening the corners, the union $\Delta_{j}\sqcup_{K_{j}}\widetilde{\Delta}_{j}$ defines a $2$-sphere, say $L_{j}$, in $V_{1}(p,q,2)$ which is clearly a suspension of $K_{j}$ (for each $j=1,...,\ell$). Moreover, since $K_{j}$’s are disjointly embedded in $V_{1}(p,q)$ so are $L_{j}$’s in $V_{1}(p,q,2)$ by construction.
Now using the same rule, one can also define the map $\Delta:\mathbb{C}^{2}\times[0,1]\to\mathbb{C}^{3}\cong\mathbb{R}^{6}$ and $\widetilde{\Delta}:\mathbb{C}^{2}\times[-1,0]\to\mathbb{C}^{3}\cong\mathbb{R}^{6}$ given by
$\Delta(x_{0}+iy_{0},x_{1}+iy_{1},t)=(f(t)^{1/p}\,x_{0},f(t)^{1/p}\,y_{0},f(t)^{1/q}\,x_{1},f(t)^{1/q}\,y_{1},g(t),0)$,
$\widetilde{\Delta}(x_{0}+iy_{0},x_{1}+iy_{1},t)=(f(t)^{1/p}\,x_{0},f(t)^{1/p}\,y_{0},f(t)^{1/q}\,x_{1},f(t)^{1/q}\,y_{1},-g(t),0)$.
The restriction of $\Delta$ to $\mathbb{C}^{2}\times(0,1)$ is one-to-one and an immersion, so the composition $\Delta_{j}=\Delta\circ\Pi_{j}$ is an embedding of $D^{2}\times(0,1)$ because $\Pi_{j}$ is an embedding. Moreover, it extends (smoothly) to an embedding $\Delta_{j}:D^{2}\times[0,1]\hookrightarrow\mathbb{C}^{3}$ of upper hemisphere of $L_{j}$ so that $\Delta_{j}|_{S^{1}\times\{0\}}$ is the embedding $\pi_{j}$ of $K_{j}$ (the equator of $L_{j}$). Similarly, the restriction of $\widetilde{\Delta}$ to $\mathbb{C}^{2}\times(-1,0)$ is one-to-one and an immersion, so the composition $\widetilde{\Delta}_{j}=\widetilde{\Delta}\circ\Pi_{j}$ is an embedding of $D^{2}\times(-1,0)$ because $\Pi_{j}$ is an embedding. It also extends (smoothly) to an embedding $\widetilde{\Delta}_{j}:D^{2}\times[-1,0]\hookrightarrow\mathbb{C}^{3}$ of lower hemisphere of $L_{j}$ so that $\widetilde{\Delta}_{j}|_{S^{1}\times\{0\}}$ is the embedding $\pi_{j}$ of $K_{j}$. Therefore, these two embeddings together form the 2-sphere $L_{j}$ whose equator is $K_{j}$.
Observe the images of $F\subset\mathbb{C}^{2}\times\{0\}$ under the restrictions $\Delta|_{\mathbb{C}^{2}\times\{0\}}$ and $\widetilde{\Delta}|_{\mathbb{C}^{2}\times\{0\}}$ satisfy
$z_{0}^{p}+z_{1}^{q}+z_{2}^{2}=1$
@@
(which defines the Brieskorn variety $\widetilde{F}:=V_{1}(p,q,2)$), and the image of $\partial F=\Sigma(p,q)$ is contained in $\partial\widetilde{F}=\Sigma(p,q,2)$. In other words, $\Delta$ and $\widetilde{\Delta}$ provide a proper embedding of $F$ into $\widetilde{F}$. Furthermore, as explained in the proof of Proposition 2.1, $\widetilde{F}$ and $F$ are the regular fibers of a compatible relative Lefschetz fibration on the Stein pair $(D^{6},D^{4})$, so $F$ is a Weinstein subdomain of $\widetilde{F}$ as claimed in (v). In particular, $\widetilde{F}$ is the page of the (induced) compatible open book on $(S^{5},\xi^{5}_{std})$, and so its boundary (the binding) $\Sigma(p,q,2)$ is a contact submanifold of $(S^{5},\xi^{5}_{std})$ and the binding $\Sigma(p,q)$ of the (induced) compatible open book on $(S^{3},\xi^{3}_{std})$ is contactly embedded in $\Sigma(p,q,2)$.
In order to check (i), first we observe the following: For any $(\theta,t)\in S^{1}\times[0,1]$ the coordinates of the image $\Delta_{j}(\theta,t)$ together satisfy the equation $z_{0}^{p}+z_{1}^{q}+z_{2}^{2}=1$. (Thanks to the property (P1)) Same equation is satisfied under $\widetilde{\Delta}_{j}$ for any $(\theta,t)\in S^{1}\times[-1,0]$. So, $L_{j}$ is lies on $\widetilde{F}$.
Now consider the restriction of $\Delta:\mathbb{C}^{2}\times[0,1]\to\mathbb{C}^{3}$ to $S^{3}\times[0,1]$. So, we have the homotopy
$\Delta|_{S^{3}\times[0,1]}:S^{3}\times[0,1]\to S^{5}$
of $3$-spheres in $S^{5}$. When $t=0$, this restriction is exactly the standard contact $3$-sphere $(S^{3},\xi^{3}_{std})$ that we started with in the first place. Denote by $S^{3}_{t}$ the image of $S^{3}\times\{t\}$ under $\Delta|_{S^{3}\times[0,1]}$. If we pull-back the standard contact form $\alpha_{std}^{5}$ on $S^{5}$, we obtain the resulting $1$-form
$\alpha_{t}=(\Delta|_{S^{3}\times\{t\}})^{*}(\alpha_{std}^{5})={f(t)^{\frac{2}{p}}}(x_{0}dy_{0}-y_{0}dx_{0})+{f(t)^{\frac{2}{q}}}(x_{1}dy_{1}-y_{1}dx_{1})$
which defines a contact form on $S^{3}_{t}$: Since the coordinates $z_{0}$ and $z_{1}$ of $S^{3}\subset S^{5}$ cannot be zero at the same time, we have
$\alpha_{t}\wedge d\alpha_{t}=2f(t)^{\frac{2}{p}+\frac{2}{q}}(x_{0}dy_{0}\wedge dx_{1}\wedge dy_{1}-y_{0}dx_{0}\wedge dx_{1}\wedge dy_{1}+x_{1}dx_{0}\wedge dy_{0}\wedge dy_{1}-y_{1}dx_{0}\wedge dy_{0}\wedge dx_{1})\neq 0.$
Indeed, for each $10\leq t<$ 1, $(S^{3}_{t},\xi_{t}=\text{Ker}(\alpha_{t}))$ is contactomorphic to $(S^{3},\xi^{3}_{std})$, (see Proposition 2.5 in [14]). The same argument applies for $\widetilde{\Delta}$ as well. Also at the level $t=0$, on both $\Delta$ and $\widetilde{\Delta}$, we have the Legendrian $K_{j}$ in $(S^{3},\xi^{3}_{std})$. As the parameter $t$ changes, we obtain $K_{j}\times\{t\}$ that is embedded in the contact submanifold $S^{3}_{t}$ for each $t$. Moreover, Proposition 2.5 in [14] implies further: We have a $1$-parameter smooth family of contact $1$-forms $\alpha_{t}$, $-1<t<1$, which gives rise to the smooth family of hyperplane distributions $\xi_{t}$ on $S^{3}_{t}$ By the above pull-back discussion (and by Proposition 2.5), one can see that each $K_{j}\times\{t\}$ is Legendrian in $S^{3}_{t}$, and so isotropic in $(S^{5},\xi^{5}_{std})$. Therefore, being the union of isotropic circles, $L_{j}$ is Legendrian in $(S^{5},\xi^{5}_{std})$. As a result, being a Legendrian on a page of the compatible open book of $(S^{5},\xi^{5}_{std})$, $L_{j}$ is Lagrangian in $V_{1}(p,q,2)$. To proof (i), it remains to show that $L_{j}$ is embedded in $V_{1}(p,q,2)$:
For each $-1<t<1$, the maps
$\Delta|_{D^{4}\times[0,1)}:D^{4}\times[0,1)\to D^{6}\subset\mathbb{C}^{3}\quad\text{and}\quad\widetilde{\Delta}|_{D^{4}\times(-1,0]}:D^{4}\times(-1,0]\to D^{6}\subset\mathbb{C}^{3}$
symplectically embeds $(D_{t}^{4}=D^{4}\times\{t\},d\alpha_{t})$ into $(D^{6},d\alpha_{std}^{5})$. Therefore, the images $D_{j}\times\{t\}$ of $D^{2}$ are embedded submanifolds of $D_{t}^{4}$ with Legendrian boundaries $K_{j}\times\{t\}$ in $(S^{3}_{t},\xi_{t})$. Indeed, each $D_{j}\times\{t\}$ of $D^{2}$ is Lagrangian in $(D_{t}^{4},d\alpha_{t})$ because of the fact that exact Lagrangian fillings are invariant under Legendrian isotopies. Therefore, (i) follows.
@@
Finally, (iii) follows from Lemma 3.2 below which applies to all (odd) dimensions. This concludes the proof of Proposition 3.1.
@@
between $\Xi_{0}=\Xi_{J}$ and $\Xi_{1}=\Xi_{\widetilde{J}}$ as claimed. ∎

### v3

@@ 1. Introduction
Let $E$ be a compact $2n$-dimensional manifold with corners equipped with an exact symplectic form $\omega=d\lambda$ such that the both faces of the boundary $\partial E=\partial_{v}E\cup\partial_{h}E$ are convex. A compatible Lefschetz fibration on a compact $2n$-dimensional manifold with corners $E$ is a holomorphic map $\pi:E^{2n}\rightarrow D^{2}$ for some choice of compatible complex structure on $E$, which satisfies the following conditions:
@@
    The map $\pi$ has finitely many nondegenerate distinct critical values $s_{1},...,s_{k}\in int(D^{2})$, and there exists an unique critical point $r_{j}\in\pi^{-1}(s_{j})$ for each $j=1,...,k$, and compatible complex structure on $E$ is integrable near the critical points,
@@
    Near each critical point $r_{j}$ and the corresponding critical value $s_{j}$, there are local complex coordinate charts matching with the orientations of $E$ and $D^{2}$ such that the map $\pi$ locally has the form $\pi(z_{1},...,z_{n})=z_{1}^{2}+...+z_{n}^{2}$,
@@
    The restriction of the map $\pi$ to $E\backslash\pi^{-1}(\{s_{1},...,s_{k}\})$ is a locally trivial fibration whose fibers are $(2n-2)$-dimensional exact symplectic manifolds with convex boundary for the restricted forms,
  4. (4)
    The restriction of the map $\pi|_{\partial_{v}E}:\partial_{v}E\rightarrow\partial D^{2}$ to the vertical boundary $\partial_{v}E=\pi^{-1}(\partial D^{2})$ is a surjective smooth fiber bundle. Moreover, there is a neighborhood of the horizontal boundary $\partial_{h}E=\bigcup\limits_{z\in D^{2}}\partial(\pi^{-1}(z))$ such that the restriction of $\pi$ to this neighborhood is a product fibration $D^{2}\times\mathcal{N}(\partial F)$ where $F$ is a regular fiber of $\pi$ and $\mathcal{N}(\partial F)$ is a neighborhood of $\partial F$. The restricted $\omega$ and $\lambda$ to this neighborhood are sum of forms from the two factors. Also, $\pi|_{\partial_{h}E}:\partial_{h}E\rightarrow D^{2}$ are surjective fiber bundles.
The corners of $E$ can be rounded to get a Liouville domain so that one can obtain an exact symplectic Lefschetz fibration on the Liouville domain over disk whose regular fibers are Liouville subdomains equipped with the restriction of $\lambda$. If $E$ has a Weinstein structure endowed with a Lefschetz fibration whose fibers are $(2n-2)$-dimensional Weinstein subdomains, then by rounding the corners of $E$ one can get a compatible Lefschetz fibration $\pi:W\rightarrow D^{2}$ on the Weinstein domain $W$ such that each regular fiber is a Weinstein subdomain with the restricted structure.
@@
A pair $(W^{6},X^{4})$ is a relative Stein pair if $W$ is a Stein domain, $X$ is a Stein subdomain of $W$, and such that the induced pair $(W,X)$ of symplectic manifolds admits a Weinstein structure in which the handles are relative (relative handle decomposition) in the following sense:
@@
The main idea of this definition is the following: While we create the Stein domain $W$ handle by handle, the Stein domain $X$ is evolving as a subdomain of $W$ in the meantime. The Liouville vector field of $W$ is hence tangent to $X$, with all critical points contained in $X$, and the index of the critical point in $X$ being less than or equal to the index of the critical point in $W$. Moreover, the advantage of the conditions in Definition 1.1 is that “admissible Stein pairs” (see Definition 1.8), which form a subclass of relative Stein pairs, can be described by means of special diagrams which we call relative Stein diagrams. Such diagrams are, indeed, relative versions of standard Stein diagrams of Stein surfaces introduced by Gompf in [11].
@@
where $\widetilde{W}$ is a subcritical Stein $6$-domain which splits as $\widetilde{W}=\widetilde{X}\times D^{2}$, and for each $j=1,...,\ell$ the pair $(H_{j},h_{j})$ is a relative $3$-handle attached along a suspendible Legendrian $2$-link.
@@
Suspendibility of $2$-links will be introduced in Definition 3.4 and Definition 3.6. For an admissible Stein pair $(W,X)$, its subcritical part $(\widetilde{W},\widetilde{X})$ satisfies the conditions $(0)$, $(1)$ and $(2)$ of Definition 1.1 (see the proof of Proposition 4.3 or Remark 3.5). Also by definition of admissibility, the condition $(3)$ of Definition 1.1 is obviously satisfied. Hence, after attaching relative $3$-handles to the relative Stein pair $(\widetilde{W},\widetilde{X})$ we obtain a new relative Stein pair $(W,X)$. Therefore, admissible Stein pairs, indeed, form a special subclass of relative Stein pairs.
@@ 2. BRIESKORN VARIETIES AS SEIFERT FILLINGS
By considering the coordinates $(z_{0},z_{1},z_{2})$ on $\mathbb{C}^{3}$, we extend $\pi$ to obtain Lefschetz fibration
@@
The map $\widetilde{\pi}$ is injective on the set of its $(p-1)(q-1)$ nondegenerate distinct critical points, so it has $(p-1)(q-1)$ vanishing cycles corresponding to its critical values. The regular fiber $\widetilde{F}=\widetilde{\pi}^{-1}(\epsilon)$ of $\widetilde{\pi}$ is the Brieskorn variety $V_{\epsilon}(p,q,2)$. Note that $\pi$ and $\widetilde{\pi}$ have the same morsification term and if we intersect the fiber $\widetilde{F}$ with $z_{2}=0$ plane, we obtain the fiber $F$. Also, we have $F\setminus\partial F\subset\widetilde{F}\setminus\partial\widetilde{F}$ and $\partial F=\partial\widetilde{F}\cap F$ which means that the regular fibers of $\pi$ is properly embedded in those of $\widetilde{\pi}$. Moreover, as discussed above the critical points of $\widetilde{\pi}$ are the same as those of $\pi$ (with additional $z_{2}$-coordinates). To understand the vanishing cycles of $\widetilde{\pi}$, one can investigate the Lefschetz fibration
$\Pi:\widetilde{F}\to D^{2},\quad\Pi(z_{0},z_{1},z_{2})=z_{2}$
on the fiber $\widetilde{F}$. For each critical point of $\pi$, there are exactly two critical points of $\Pi$. So, for each vanishing cycle in $\pi^{-1}(0)$, there are two vanishing cycles in the fiber of $\Pi$. Since these two vanishing cycles coincide, one can form a matching sphere in ${\widetilde{F}}$ by joining the corresponding two critical points of $\Pi$. These matching spheres form a basis of vanishing cycles in $\widetilde{F}$ for the fibration $\widetilde{\pi}$. They can be taken to intersect in the fiber $\widetilde{\pi}^{-1}(0)$ according to the same graph as the vanishing cycles in $\pi^{-1}(0)$, (see [19]). In conclusion, the vanishing cycles in the fiber $\widetilde{F}$ of $\widetilde{\pi}$ are these Lagrangian spheres which we call the suspensions of those in the fiber $F$ of $\pi$. Hence, introducing a $z_{2}^{2}$ term extends (“suspends”) the Lefschetz fibration $\pi$ on $D^{4}$ to the Lefschetz fibration $\widetilde{\pi}$ on $D^{6}$, and the monodromy $\widetilde{\varphi}_{p,q}$ of $\widetilde{\pi}$ is the monodromy of the fibered Brieskorn manifold $\Sigma(p,q,2)\subset S^{5}$ which is the product of $(p-1)(q-1)$ right-handed Dehn twists along Lagrangian spheres which are the suspensions of the corresponding vanishing cycles of $\pi$. In particular, note that the binding $\Sigma(p,q)$ of the induced open book on $S^{3}$ is properly embedded in the binding $\Sigma(p,q,2)$ of the induced open book on $S^{5}$.
@@
we see that $(-1/27,0)$ and $(1/27,0)$ are the critical points of $\pi$. By substituting these into $\pi$, we get the corresponding critical values $-0.00020322105$ and $-0.00010161052$, respectively.
Then one needs to calculate the determinant of the Hessian matrix at the critical points. The Hessian matrix is given by (see [7] for a complete discussion):
@@
Figure 4. The regular fiber and the monodromy curves of the torus knot $\Sigma(3,2)$ (trefoil).
@@ 3. Suspended Legendrian 2-links in the complement of Brieskorn manifolds
By [18], we may assume there exists a torus knot $\Sigma(p,q)$ such that the equatorial Legendrian $1$-link $K\subset(S^{3},\xi^{3}_{std})$ is embedded in the interior of a Seifert surface $F:=V_{\epsilon}(p,q)$ of $\Sigma(p,q)$. Also, since $F$ is the page of a compatible open book (with binding $\Sigma(p,q)$) on $(S^{3},\xi^{3}_{std})$, the framing of each component of $K$ in $F$ is equal to its contact framing in $(S^{3},\xi^{3}_{std})$ (see [1]). Therefore, (ii) and (iv) follow. By assumption, for each $j=1,...,\ell$, there exists an embedded Lagrangian disk $D_{j}$ in $D^{4}\subset\mathbb{C}^{2}$ such that the boundary is the embedded Legendrian $K_{j}$ in $S^{3}$. In what follows, we construct the Lagrangian $2$-knot $L_{j}$ in the Brieskorn variety $V_{\epsilon}(p,q,2)$ by lifting the disk $D_{j}$ under a double branched cover. We will also make use of the fact that $V_{\epsilon}(p,q,2)$ is obtained by suspending the Seifert surface $F=V_{\epsilon}(p,q)$ of $K$ as observed in the proof of Proposition 2.1. Consider the map
$f:V_{\epsilon}(p,q,2)\to\mathbb{C}^{2},\quad f(z_{0},z_{1},z_{2})=(z_{0},z_{1})$
where $(z_{0},z_{1},z_{2})$ are complex coordinates on $\mathbb{C}^{3}$. This map is a two-fold cover branched along the fiber $\pi^{-1}(0)$ in $\mathbb{C}^{2}$ where $\pi$ is the Lefschetz fibration given in the proof of Proposition 2.1. For every point except on the fiber $\pi^{-1}(0)$, there are two preimage points of $f$ which are related by changing the sign on the $z_{2}$-coordinate. A disk that is not intersecting the fiber $\pi^{-1}(0)$ lifts to two disjoint embedded disks in $V_{\epsilon}(p,q,2)$. Since $f$ is a holomorphic covering map away from $\pi^{-1}(0)$, we may assume that $f$ is a symplectomorphism so that the two lifted disks are Lagrangian. If the disk has boundary equal to $K_{j}$ in $\pi^{-1}(0)$ and a neighborhood of the boundary of the disk projects to a line $[0,\epsilon)$ in $\mathbb{C}$ under $\pi$, then the two lifts of $D_{j}$ can be glued to form a smooth Lagrangian sphere in $V_{\epsilon}(p,q,2)$ whose equator projects to $(-\epsilon,\epsilon)$ under the Lefschetz fibration $\Pi$ on $V_{\epsilon}(p,q,2)$ also given in the proof of Proposition 2.1. Note that, if you take $(z^{2}\circ\Pi)$ or $(\Pi)^{2}$ is equal to $\pi\circ f$. Therefore, (i) follows.
@@
As explained in the proof of Proposition 2.1, $\widetilde{F}=V_{\epsilon}(p,q,2)$ and $F$ are the regular fibers of a compatible relative Lefschetz fibration on the Stein pair $(D^{6},D^{4})$, so $F$ is a Weinstein subdomain of $\widetilde{F}$ as claimed in (v). In particular, $\widetilde{F}$ is the page of the (induced) compatible open book on $(S^{5},\xi^{5}_{std})$, and so its boundary (the binding) $\Sigma(p,q,2)$ is a contact submanifold of $(S^{5},\xi^{5}_{std})$ and the binding $\Sigma(p,q)$ of the (induced) compatible open book on $(S^{3},\xi^{3}_{std})$ is contactly embedded in $\Sigma(p,q,2)$.
@@
Finally, (iii) follows from Lemma 3.2 below which applies to all (odd) dimensions. This concludes the proof of Proposition 3.1. ∎
@@
between $\Xi_{0}=\Xi_{J}$ and $\Xi_{1}=\Xi_{\widetilde{J}}$ as claimed.